Calculate the mass of sulphuric(VI) acid 250cm3 of a solution whose concentration is 0.25mole dm-3
1
Expert's answer
2021-09-10T00:40:35-0400

Solution:

sulphuric(VI) acid = H2SO4


Convert cm3 to dm3:

1 cm3 = 0.001 dm3

Hence,

(250 cm3) × (0.001 dm3 / 1 cm3) = 0.25 dm3


Calculate the moles of H2SO4:

(0.25 mol dm-3) × (0.25 dm3) = 0.0625 mol H2SO4


Calculate the mass of H2SO4:

The molar mass of H2SO4 is 98.08 g/mol

Hence,

(0.0625 mol H2SO4) × (98.08 g H2SO4 / 1 mol H2SO4) = 6.13 g H2SO4


Answer: The mass of sulphuric(VI) acid (H2SO4) is 6.13 g